OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-7, 6) and is perpendicular to -7x + 9y = -2 Please show all of your work.

OpenStudy (anonymous):

-7x + 9y = -2 9y = 7x - 2 y= 7/9x - 2/9 perpendicular meaning m1 x m2 = -1 m1 = 7/9 so m2 = -9/7 now using the point (-7,6) sub it into the equation y - y1 = m (x - x1) m= use the m2 gradient y1 and x1 = the point so therefore y-(6) = -9/7 (x--7) y-6 = -9/7 (x+7) y-6 = -9/7x -9+6 y-6 = -9/7x -3